OpenStudy (anonymous):

The peregrine falcon is the fastest bird flying at a speed of 200mi/h. Nature has adapted the bird to reach such a speed by placing baffles in its nose to prevent air from rushing in and slowing it down. Also, the bird's eyes adjust their focus faster than the eyes of any other creature, so the falcon can focus quickly on its prey.Assume that a peregrine falcon is moving horizontally at its top speed at a height of 100 m above the ground when it brings it swing into its sides and begins to drop in free fall. How far will the bird fall vertically while traveling horizontally a distance of100 m

OpenStudy (anonymous):

First thing you need to do is convert 200 miles per hour (mi/h) to meters per second (m/s). Then draw a diagram showing the horizontal and vertical components of the free fall. Horizontal being ( 200 mi/h converted to m/s ) and the vertical component being gravity ( -9.8 m/s ). Use a kinematics equation to find the time it takes to travel 120 m horizontally and apply that time to the vertical component to find its distance.

OpenStudy (anonymous):

At a horizontal top speed of 200 mph, the time it will take the falcon to travel a horizontal distance of 80 m would be T = 80/(200 * 0.44704) NOTE the factor "0.44704" converts "mi/hour" to "m/sec" T = 0.895 sec. Next working formula is Y = Vy(T) + (1/2)gT^2 where Y = vertical distance that the bird will make Vy = vertical component of the bird's velocity = 0 T = time interval = 0.895 sec (as calculated above) g = acceleration due to gravity = 9.8 m/sec^2 Substituting values, Y = 0 + (1/2)(9.8)(0.895)^2 Y = 3.92 meters